# local non-bossiness

### v3

@@ A $\mu$-improving cycle is a tuple of different students $(i_{1},\ldots,i_{r})$ such that
    A $\mu$-improving cycle $(i_{1},\ldots,i_{r})$ is implementable by $\mu^{\prime}$ when
    $\quad\quad\mu^{\prime}(i_{l})=\mu(i_{l+1}),\quad\,\,\forall l\in\{1,\ldots,r\}\,\,{\mbox{[modulo $r$]}}.$
@@ A $\mu$-improving cycle $(i_{1},\ldots,i_{r})$ is implementable by $\mu^{\prime}$ when
    Each student $i\in I$ is part of a $\mu$-improving cycle implementable by $\mu^{\prime}$ and contained in the graph $G$ (see Lemma 5 in Appendix B).
@@
    Any student that $\mu$-blocks an edge in $E$ belongs to $V$.
    If $k\in N$ $\mu$-blocks an edge $[i,j]$, then $\mu(j)P_{k}\mu(k)$ and $k\succ_{\mu(j)}i$. Since $[i,j]\in E$ implies that $\mu(j)=\mu^{\prime}(i)$, we have that $\mu^{\prime}(i)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(i)}i$. Hence, there are two possibilities:
    • $\bullet$
      If $k\neq 1$, the stability of $\mu^{\prime}$ implies that $\mu^{\prime}(k)R_{k}\mu^{\prime}(i)$. Since $\mu^{\prime}(i)P_{k}\mu(k)$, the transitivity of $P_{k}$ ensures that $k\in I$. Hence, $k\in V$.
    • $\bullet$
      If $k=1$, the previous argument cannot be applied, because student $1$ has different preferences in $(P_{1},P_{-1})$ and $(P^{\prime}_{1},P_{-1})$. However, the property (i) ensures that $i_{1}$ is part of a $\mu$-improving cycle implementable by $\mu^{\prime}$ that is contained in $G$. Hence, there is $i\in I$ such that $[i,i_{1}]\in E$. We conclude that $1\in V$ because there is an edge between $i$ and every student assigned to school $\mu(i_{1})=\overline{s}=\mu(1)$.
  • (iii)
@@
    By construction, each node in $V\setminus I$ has a positive in-degree. Moreover, the property (i) ensures that each student $i\in I$ is part of a $\mu$-improving cycle implementable by $\mu^{\prime}$ that is contained in $G$. Hence, there exists $j\in V$ such that $[j,i]\in E$.
@@
Let $G^{\prime}=(V,E^{\prime})$ be the graph obtained from $G$ through the following edge-replacement procedure: given a student $j\in N$, substitute each $[i,j]\in E$ for which the set $I[i,j]\equiv\{k\in N:\,\mbox{$k$ $\mu$-blocks $[i,j]$}\}$ is non-empty by the directed edge $[\overline{k},j]$, where $\overline{k}$ is the student with the highest priority at school $\mu(j)$ among those in $I[i,j]$. Notice that, when $[\overline{k},j]\notin E$, its inclusion is feasible because $\overline{k}\in V$ (see property (ii)).
It follows that $G^{\prime}$ has no $\mu$-improving cycles. Indeed, as any student that $\mu$-blocks an edge in $E$ belongs to $V$, [^10] Since the student $1$ may $\mu$-block edges in $E$, it is crucial to have $1\in V$. The existence of $i_{1}\in\mu(\overline{s})$ such that $\mu^{\prime}(i_{1})\neq\overline{s}$ is only used to prove it (see the property (ii) for details). the construction of the new set of edges $E^{\prime}$ guarantees that a $\mu$-improving cycle in $G^{\prime}$ cannot be blocked by any student. Hence, if there is such a cycle, then we can implement it contradicting the fact that $\mu$ is the student-optimal stable matching for $(P_{1},P_{-1})$.
@@
However, property (iii) ensures that all vertices in $V$ have a positive in-degree in $G^{\prime}$. Indeed, when $E^{\prime}$ is constructed from $E$, if some directed edges of the form $[i,j]$ are deleted, then an edge who points to $j$ is included. As a consequence, $G^{\prime}$ has a cycle. Since any cycle in $G^{\prime}$ is a $\mu$-improving cycle, we obtain a contradiction. [^11] If a student $k$ $\mu$-blocks $[i,j]\in E$, then $\mu(j)P_{k}\mu(k)$. Hence, each $[\overline{k},j]\in E^{\prime}\setminus E$ satisfies $\mu(j)P_{\overline{k}}\mu(\overline{k})$. For this reason, any cycle in $G^{\prime}$ is a $\mu$-improving cycle. However, a $\mu$-improving cycle in $G^{\prime}$ is not necessarily implemented by $\mu^{\prime}$.
@@
As Theorem 1 demonstrates, when the ${\rm{DA}}$ mechanism is implemented, no student can modify her preferences to change her classmates without changing her school. But what happens when multiple students, all assigned to the same school, change their preferences? Can they change their classmates without changing their school? The answer is negative, and not only for ${\rm{DA}}$, but also for any locally non-bossy and strategy-proof mechanism (see Lemma 2).
@@ 4. A characterization of DA via local non-bossiness
The proof is given in the Appendix B.
@@
we have that $\Omega(\overline{N},P)={\rm{DA}}^{\succ}(\overline{N},P)=((1,s_{2}),(2,s_{1}),(3,s_{1}),(4,s_{0})).$ If $P^{\prime}_{2}:s_{2},s_{1},s_{0}$, $\Omega(\overline{N},(P^{\prime}_{2},P_{-2}))={\rm{DA}}^{\succ^{\prime}}(\overline{N},(P^{\prime}_{2},P_{-2}))=((1,s_{2}),(2,s_{1}),(3,s_{0}),(4,s_{1})).$ Therefore, the student $2$ can change her classmates without modify her school. [^13] This argument also shows that $\Omega$ is not group strategy-proof. Indeed, the coalition $\{2,4\}$ can report preferences $(P^{\prime}_{2},P_{4})$ to improve the assignment of $4$ without hurting $2$. Hence, the same example allows us to show that group strategy-proofness is a critical axiom in Corollary 2. $\Box$
@@ 5. Mechanisms under preferences over colleagues
Most of the literature in school choice assumes that students only care about the school to which they are assigned. However, it is unrealistic that students do not care about the assignment of others. In particular, it has been empirically documented that the identity of classmates is an important consideration when students are choosing a school (Rothstein, 2006; Abdulkadiroğlu et al., 2020; Allende, 2021; Che et al., 2022; Beuermann et al., 2023; Cox et al., 2023).
When students have arbitrary preferences over the set of matchings, many of the results in the literature break down. Specifically, a stable matching may not exist, even when students only care about the identity of the other students assigned to her school (Echenique and Yenmez, 2007). One way to overcome this problem is to restrict the domain of preferences. For instance, assuming that students have school-lexicographic preferences, meaning that they are primarily concerned with the school to which they are assigned (Sasaki and Toda, 1996; Dutta and Massó, 1997). In this domain, some of the standard results are recovered, such as the existence of a stable matching. However, a stable and strategy proof mechanism may not exist (Duque and Torres-Martínez, 2023). One may wonder if there is a subdomain of school-lexicographic preferences where stability and strategy-proofness are compatible. In this section, we show that within the subdomain of school-lexicographic preferences where only the classmates matter, the local non-bossiness of ${\rm{DA}}$ implies that this rule induces a stable and strategy-proof mechanism.
@@
Within the domain of school-lexicographic preferences, let us consider a scenario where students are solely concerned about the composition of the school to which they are assigned. Consequently, a student will be indifferent between two matchings when she is assigned to the same school with the same classmates. This is the idea of the following subdomain of preferences.
@@
In words, when $(\unrhd_{i})_{i\in N}$ belongs to $\mathcal{D}_{{\rm{c}}}$, each student $i$ strictly prefers $\mu$ to $\eta$ only when she is assigned to different schools or to the same school with different classmates.
@@ 6. Concluding remarks
In this paper, we have shown that the bossiness of the ${\rm{DA}}$ mechanism is limited. When ${\rm{DA}}$ is implemented, it is well known that a student can affect another student’s assignment by changing her preferences without modifying her school. We have demonstrated that this holds only for students not assigned to the same school. In other words, under ${\rm{DA}}$, a student cannot change her classmates without changing her own school. Additionally, for any strategy-proof mechanism, local non-bossiness guarantees that no coalition of students assigned to the same school can misrepresent their preferences to either improve their assignments or maintain their school while modifying their classmates. Since ${\rm{DA}}$ is not group strategy-proof, our result implies that a successful manipulating coalition of DA must include students from different schools.
Local non-bossiness is not only interesting in itself but also because of its application to school choice problems with externalities. In this framework, even when students prioritize their assigned school first and then consider the assignment of others, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). However, if we restrict preferences to be such that only the school and the classmates matter, it turns out that ${\rm{DA}}$ induces the only stable and strategy-proof mechanism. The crucial property behind this result is precisely its local non-bossiness.
@@
Hence, local group non-bossiness ensures that no coalition of students assigned to the same school can change the assignment of some of their classmates by reporting different preferences, without changing their school.
@@
We claim that $\Omega$ is locally bossy. Let $P_{1}$ be such that ${P}_{1}:s_{2},s_{1},s_{0}$. It is easy to see that $\Omega(P_{1},\overline{P}_{-1})={\rm{DA}}(P_{1},\overline{P}_{-1})=((1,s_{1}),(2,s_{2}),(3,s_{1}))$. Since $\Omega(\overline{P}_{1},\overline{P}_{-1})={\rm{DA}}^{S}(\overline{P})$, when student $1$’s preferences change from $P_{1}$ to $\overline{P}_{1}$, she remains assigned to school $s_{1}$, but her classmates change as $\Omega_{s_{1}}(\overline{P}_{1},\overline{P}_{-1})=\{1,2\}\neq\{1,3\}=\Omega_{s_{1}}(P_{1},\overline{P}_{-1}).$
@@
We have that $\Omega(P)=((1,s_{1}),(2,s_{1}),(3,s_{2}))$ and $\Omega_{s_{1}}(P)=\{1,2\}$. If the student $1$ changes her preference to $P^{\prime}_{1}$ such that $P^{\prime}_{1}:s_{1},s_{2},s_{0}$, she remains assigned to school $s_{1}$ but is left without classmates because $\Omega_{s_{1}}(P^{\prime}_{1},P_{2},P_{3})=\{1\}$. This argument also shows that $\Omega$ is weak bossy, because $\Omega_{s_{0}}(P)=\emptyset$ and $\Omega_{s_{0}}(P^{\prime}_{1},P_{2},P_{3})=\{2,3\}$. $\Box$
@@
It is easy to verify that $\Omega$ is strategy-proof. We claim that $\Omega$ is not locally group strategy-proof. Let $P=(P_{1},P_{2},P_{3})$ be a preference profile such that $P_{1}:s_{1},s_{2},s_{0}$ and $P_{2}:s_{2},s_{1},s_{0}$. Let $P^{\prime}_{1},P^{\prime}_{2}\in{\mathcal{L}}$ be such that $P^{\prime}_{1}:s_{1},s_{0},s_{2}$ and $P^{\prime}_{2}=P_{2}$. It follows that the coalition of classmates $\{1,2\}\in\Omega_{s_{1}}(P)$ can manipulate the mechanism $\Omega$ at $P$, because $\Omega_{1}(P)=s_{1}=\Omega_{1}(P^{\prime}_{1},P^{\prime}_{2},P_{3})$ and $\Omega_{2}(P^{\prime}_{1},P^{\prime}_{2},P_{3})=s_{2}P_{2}s_{1}=\Omega_{2}(P)$. $\Box$
@@
If $\mu^{\prime}(i)P_{i}\mu(i)$, then $i$ is part of a $\mu$-improving cycle implementable by $\mu^{\prime}$.
@@
As a consequence, there is a cycle of schools $(s_{1},\ldots,s_{r})$ in $G^{*}$ such that $(s_{1},s_{2})=(a,a^{\prime})$. Indeed, we can form this cycle departing from node $a$ in direction to node $a^{\prime}$ and moving around the graph, because after enter a node $s_{j}\neq s_{1}$ from an edge, we are always able to exit from another edge. Since $V^{*}$ is a finite set, we will eventually return to $s_{1}$. The construction of $E^{*}$ allow us to associated to $(s_{1},\ldots,s_{r})$ a group of students $i_{1},\ldots,i_{r}$ such that $i_{1}=i$, $\mu(i_{l})=s_{l}$ for all $l\in\{1,\ldots,r\}$, and $\mu^{\prime}(i_{l})=s_{l+1}$ for all $l\in\{1,\ldots,r\}$ [modulo $r$]. In particular, although the schools $s_{1},\ldots,s_{r}$ are not necessarily different, we can assume that the students $i_{1},\ldots,i_{r}$ are.
Since $\mu^{\prime}(i_{l})=s_{l+1}$, when $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ the stability of $\mu$ ensures that $j\succ_{s_{l+1}}i_{l}$ for all $j\in\mu(s_{l+1})$. Hence, the stability of $\mu^{\prime}$ implies that any edge between $s_{l+1}$ and $s_{l+2}$ is associated to a student that improves we she moves from $\mu$ to $\mu^{\prime}$. In particular, $\mu^{\prime}(i_{l+1})P_{i_{l+1}}\mu(i_{l+1})$. Therefore, $\mu^{\prime}(i_{1})P_{i_{1}}\mu(i_{1})$ ensures that $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ for all $l\in\{1,\ldots,r\}$. We conclude that $(i_{1},\ldots,i_{r})$ is a $\mu$-improving cycle implementable by $\mu^{\prime}$ that includes the student $i_{1}=i$.∎

### v4

@@ A $\mu$-improving cycle is a tuple of different students $(i_{1},\ldots,i_{r})$ such that
    Each $i\in I$ is part of a $\mu$-improving cycle contained in $G$ (see Lemma 5 in Appendix B).
@@
    By construction, each node in $V\setminus I$ has a positive in-degree. Moreover, the property (i) ensures that each student $i\in I$ is part of a $\mu$-improving cycle contained in $G$. Hence, there exists $j\in V$ such that $[j,i]\in E$.
Let $G^{\prime}=(V,E^{\prime})$ be the multigraph obtained from $G$ through the following edge-replacement procedure: given a student $j\in N$, substitute each edge $[i,j]\in E$ for which
$I[i,j]\equiv\{k\in V:\,\mbox{$k$ $\mu$-blocks $[i,j]$}\}\neq\emptyset$
@@
by the directed edge $[\overline{k},j]$, where $\overline{k}$ is the student with the highest priority at school $\mu(j)$ among those in $I[i,j]$. Notice that, when $[\overline{k},j]\notin E$, its inclusion in $E^{\prime}$ is feasible because $\overline{k}\in I[i,j]\subseteq V$.
We make the following remarks regarding graph $G^{\prime}$.
  1. (1)
    The graph $G^{\prime}$ has a cycle $(i^{*}_{1},\ldots,i^{*}_{r})$.
@@
    It follows from (ii) that all vertices in $V$ have a positive in-degree in $G^{\prime}$. Indeed, when $E^{\prime}$ is constructed from $E$, if some directed edge $[i,j]$ is deleted, then an edge who points to $j$ is included. As a consequence, $G^{\prime}$ has a cycle $(i^{*}_{1},\ldots,i^{*}_{r})$.
  2. (2)
    Any cycle in $G^{\prime}$ is a $\mu$-improving cycle.
    Remember that, if a student $k$ $\mu$-blocks $[i,j]\in E$, then $\mu(j)P_{k}\mu(k)$. Hence, each $[k,j]\in E^{\prime}\setminus E$ satisfies $\mu(j)P_{k}\mu(k)$. Since the definition of $E$ ensure that any $[k,j]\in E^{\prime}\cap E$ satisfies $\mu(j)P_{k}\mu(k)$, we conclude that any cycle in $G^{\prime}$ is a $\mu$-improving cycle.
  3. (3)
    Any cycle in $G^{\prime}$ cannot be blocked by any student in $V$.
    It is a direct consequence of the construction of the set of edges $E^{\prime}$.
  4. (4)
    The $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by student $1$.
    Since $\mu$ is the student-optimal stable matching under $(P_{1},P_{-1})$, the $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by some student in $N\setminus V$.
    For each $[i,j]\in E^{\prime}$, we have that $[i,j]\in E$ or $[i,j]\in E^{\prime}\setminus E$. Hence, $[i,j]\in E$ or there exists $[z,j]\in E$ that is $\mu$-blocked by $i$. This implies that either $\mu(j)=\mu^{\prime}(i)$ or there exists $z\in I$ such that [$\mu(j)=\mu^{\prime}(z)$, $\mu(j)P_{i}\mu(i)$, and $i\succ_{\mu(j)}z$]. As a consequence, when a student $k$ $\mu$-blocks an edge $[i,j]\in E^{\prime}$, we have that either [$\mu^{\prime}(i)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(i)}i$] or [$\mu^{\prime}(z)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(z)}z$] .
    We claim that no student $k\neq 1$ can block an edge in $E^{\prime}$. By contradiction, suppose that $k\neq 1$ $\mu$-blocks an edge $[i,j]\in E^{\prime}$. It follows from the arguments above and the stability of $\mu^{\prime}$ that there are two possible scenarios:
      If $\mu^{\prime}(i)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(i)}i$, then $\mu^{\prime}(k)R_{k}\mu^{\prime}(i)$. Since $\mu^{\prime}(i)P_{k}\mu(k)$, the transitivity of $P_{k}$ ensures that $k\in I$. Hence, $k\in V$, which contradicts the property (3).
      If $\mu^{\prime}(z)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(z)}z$, then $\mu^{\prime}(k)R_{k}\mu^{\prime}(z)$. Since $\mu^{\prime}(z)P_{k}\mu(k)$, the transitivity of $P_{k}$ ensures that $k\in I$. Hence, $k\in V$, which contradicts the property (3).
    We conclude that $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by student $1$.
We finish the proof by showing that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$.
Since $(i^{*}_{1},\ldots,i^{*}_{r})$ is blocked by student $1$, we have that $1\notin V$. Thus, the school $\overline{s}$ does not receive in $\mu^{\prime}$ students who prefer $\mu^{\prime}$ to $\mu$.
We claim that $\mu^{\prime}(\overline{s})\subseteq\mu(\overline{s})$ or $\mu(\overline{s})\subseteq\mu^{\prime}(\overline{s})$. Indeed, there are two possible scenarios:
    If no one prefers $\mu$ to $\mu^{\prime}$, then everyone in $\mu^{\prime}(\overline{s})$ is indifferent between $\mu^{\prime}$ to $\mu$, which implies that $\mu^{\prime}(\overline{s})\subseteq\mu(\overline{s})$.
    If at least one student prefers $\mu$ to $\mu^{\prime}$, then we can repeat all the previous arguments swapping the roles of $(P_{1},P_{-1})$ and $(P^{\prime}_{1},P_{-1})$ in order to show that $\overline{s}$ does not receive in $\mu$ students who prefers $\mu$ to $\mu^{\prime}$. Notice that, if there exists $i\in\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$, then $i$ belongs to $I$. Hence, the property (i) guarantees that $i$ is part of a $\mu$-improving cycle contained in $G$, which ensures that there is $k\in I$ such that $[k,i]\in E$. This implies that $1\in V$, because $\mu^{\prime}(k)=\mu(i)=\overline{s}=\mu(1)$. A contradiction. Hence, $\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$ is an empty set, which implies that $\mu(\overline{s})\subseteq\mu^{\prime}(\overline{s})$.
Finally, since Lemma 4 guarantees that $|\mu(s)|=|\mu^{\prime}(s)|$ for all $s\in S$ such that $\mu(s)\neq\mu^{\prime}(s)$, we conclude that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$. ∎
@@
As Theorem 1 demonstrates, when the ${\rm{DA}}$ mechanism is implemented, no student can modify her preferences to change her schoolmates without changing her school. But what happens when multiple students, all assigned to the same school, change their preferences? Can they change their colleagues without changing their school? The answer is negative, and not only for ${\rm{DA}}$, but also for any locally non-bossy and strategy-proof mechanism (see Lemma 2).
@@ 4. A characterization of DA via local non-bossiness
The proof of this result and examples that show that the six axioms are independent are provided in the Appendix B.
@@
we have that $\Omega(\overline{N},P)={\rm{DA}}^{\succ}(\overline{N},P)=((1,s_{2}),(2,s_{1}),(3,s_{1}),(4,s_{0})).$ If $P^{\prime}_{2}:s_{2},s_{1},s_{0}$, $\Omega(\overline{N},(P^{\prime}_{2},P_{-2}))={\rm{DA}}^{\succ^{\prime}}(\overline{N},(P^{\prime}_{2},P_{-2}))=((1,s_{2}),(2,s_{1}),(3,s_{0}),(4,s_{1})).$ Therefore, the student $2$ can change her colleagues without modify her school. [^14] This argument also shows that $\Omega$ is not group strategy-proof. Indeed, the coalition $\{2,4\}$ can report preferences $(P^{\prime}_{2},P_{4})$ to improve the assignment of $4$ without hurting $2$. Hence, the same example allows us to show that group strategy-proofness is a critical axiom in Corollary 2. $\Box$
@@ 5. Mechanisms under preferences over colleagues
Most of the literature in school choice assumes that students only care about the school to which they are assigned. However, it is unrealistic that students do not care about the assignment of others. In particular, it has been empirically documented that the identity of colleagues is an important consideration when students are choosing a school (Rothstein, 2006; Abdulkadiroğlu et al., 2020; Allende, 2021; Che et al., 2022; Beuermann et al., 2023; Cox et al., 2023).
When students have arbitrary preferences over the set of matchings, many of the results in the literature break down. Specifically, a stable matching may not exist, even when students only care about the identity of the other students assigned to her school (Echenique and Yenmez, 2007). One way to overcome this problem is to restrict the domain of preferences. For instance, assuming that students have school-lexicographic preferences, meaning that they are primarily concerned with the school to which they are assigned (Sasaki and Toda, 1996; Dutta and Massó, 1997). In this domain, some of the standard results are recovered, such as the existence of a stable matching. However, a stable and strategy proof mechanism may not exist (Duque and Torres-Martínez, 2023). One may wonder if there is a subdomain of school-lexicographic preferences where stability and strategy-proofness are compatible. In this section, we show that within the subdomain of school-lexicographic preferences where only the schoolmates matter, the local non-bossiness of ${\rm{DA}}$ implies that this rule induces a stable and strategy-proof mechanism.
@@
Within the domain of school-lexicographic preferences, let us consider a scenario where students are solely concerned about the composition of the school to which they are assigned. Consequently, a student will be indifferent between two matchings when she is assigned to the same school with the same colleagues. This is the idea of the following subdomain of preferences.
@@
In words, when $(\unrhd_{i})_{i\in N}$ belongs to $\mathcal{D}_{{\rm{c}}}$, each student $i$ strictly prefers $\mu$ to $\eta$ only when she is assigned to different schools or to the same school with different colleagues.
@@ 6. Concluding remarks
In this paper, we have shown that the bossiness of the ${\rm{DA}}$ mechanism is limited. When ${\rm{DA}}$ is implemented, it is well known that a student can affect another student’s assignment by changing her preferences without modifying her school. We have demonstrated that this holds only for students not assigned to the same school. In other words, under ${\rm{DA}}$, a student cannot change her schoolmates without changing her own school. Additionally, for any strategy-proof mechanism, local non-bossiness guarantees that no coalition of students assigned to the same school can misrepresent their preferences to either improve their assignments or maintain their school while modifying their colleagues Since ${\rm{DA}}$ is not group strategy-proof, our result implies that a successful manipulating coalition of DA must include students from different schools.
Local non-bossiness is not only interesting in itself but also because of its application to school choice problems with externalities. In this framework, even when students prioritize their assigned school first and then consider the assignment of others, a stable and strategy-proof mechanism may not exist (Duque and Torres-Martínez, 2023). However, if we restrict preferences to be such that only the school and the schoolmates matter, it turns out that ${\rm{DA}}$ induces the only stable and strategy-proof mechanism. The crucial property behind this result is precisely its local non-bossiness.
@@
Hence, local group non-bossiness ensures that no coalition of students assigned to the same school can change the assignment of some of their colleagues by reporting different preferences, without changing their school.
@@
We claim that $\Omega$ is locally bossy. Let $P_{1}$ be such that ${P}_{1}:s_{2},s_{1},s_{0}$. It is easy to see that $\Omega(P_{1},\overline{P}_{-1})={\rm{DA}}(P_{1},\overline{P}_{-1})=((1,s_{1}),(2,s_{2}),(3,s_{1}))$. Since $\Omega(\overline{P}_{1},\overline{P}_{-1})={\rm{DA}}^{S}(\overline{P})$, when student $1$’s preferences change from $P_{1}$ to $\overline{P}_{1}$, she remains assigned to $s_{1}$, but her schoolmates change as $\Omega_{s_{1}}(\overline{P}_{1},\overline{P}_{-1})=\{1,2\}\neq\{1,3\}=\Omega_{s_{1}}(P_{1},\overline{P}_{-1}).$
@@
We have that $\Omega(P)=((1,s_{1}),(2,s_{1}),(3,s_{2}))$ and $\Omega_{s_{1}}(P)=\{1,2\}$. If the student $1$ changes her preference to $P^{\prime}_{1}$ such that $P^{\prime}_{1}:s_{1},s_{2},s_{0}$, she remains assigned to $s_{1}$ but is left without schoolmates because $\Omega_{s_{1}}(P^{\prime}_{1},P_{2},P_{3})=\{1\}$. This argument also shows that $\Omega$ is weak bossy, because $\Omega_{s_{0}}(P)=\emptyset$ and $\Omega_{s_{0}}(P^{\prime}_{1},P_{2},P_{3})=\{2,3\}$. $\Box$
@@
It is easy to verify that $\Omega$ is strategy-proof. We claim that $\Omega$ is not locally group strategy-proof. Let $P=(P_{1},P_{2},P_{3})$ be a preference profile such that $P_{1}:s_{1},s_{2},s_{0}$ and $P_{2}:s_{2},s_{1},s_{0}$. Let $P^{\prime}_{1},P^{\prime}_{2}\in{\mathcal{L}}$ be such that $P^{\prime}_{1}:s_{1},s_{0},s_{2}$ and $P^{\prime}_{2}=P_{2}$. It follows that the coalition of schoolmates $\{1,2\}\in\Omega_{s_{1}}(P)$ can manipulate the mechanism $\Omega$ at $P$, because $\Omega_{1}(P)=s_{1}=\Omega_{1}(P^{\prime}_{1},P^{\prime}_{2},P_{3})$ and $\Omega_{2}(P^{\prime}_{1},P^{\prime}_{2},P_{3})=s_{2}P_{2}s_{1}=\Omega_{2}(P)$. $\Box$
Example 7.
The school-median stable matching is locally bossy.
Given a school choice problem $[N,S,\succ,q,P]$ with $k$ stable matchings, we refer to $\mu$ as the school-median stable matching when the following properties hold:
  • If $k$ is odd, each student is assigned to her $(k+1)/2$-th (weakly) best match among all $k$ stable matchings.
  • If $k$ is even, each student is assigned to her $(k+2)/2$-th (weakly) best match among all $k$ stable matchings.
Klaus and Klijn (2006, Theorem 3.2) ensure that this matching always exists.
We claim that the mechanism that associates to each preference profile the school-median stable matching is locally bossy. Let $N=\{1,2,3,4\}$, $S=\{s_{1},s_{2}\}$, $q_{s_{1}}=q_{s_{2}}=2$, $\succ_{s_{1}}:2,1,3,4$, and $\succ_{s_{2}}:3,4,2,1.$ If preferences are given by
$P_{1}:2,1,\quad\quad P_{2}:2,1,\quad\quad P_{3}:1,2,\quad\quad P_{4}=1,2,$
there are three stable matchings:
$\mu$ $=$ $((1,s_{2}),(2,s_{2}),(3,s_{1}),(4,s_{1})),$
$\eta$ $=$ $((1,s_{1}),(2,s_{2}),(3,s_{1}),(4,s_{2})),$
$\rho$ $=$ $((1,s_{1}),(2,s_{1}),(3,s_{2}),(4,s_{2})).$
It follows that $\eta$ is the school-median stable matching. Now, if preferences of student $1$ change to $P^{\prime}_{1}:1,2$, the only stable matchings are $\eta$ and $\rho$. In this case, $\rho$ is the school-median stable matching. Therefore, as $\eta(1)=\rho(1)=s_{1}$ and $\eta(s_{1})\neq\rho(s_{1})$, we conclude that the school-median stable mechanism is locally bossy. $\Box$
@@
If $\mu^{\prime}(i)P_{i}\mu(i)$, then there exists a cycle $(i_{1},\ldots.i_{r})$ that includes $i$ and satisfies:
  • (i)
    $\mu^{\prime}(i_{l})=\mu(i_{l+1})$ for all $l\in\{1,\ldots,r\}$ [modulo $r$].
  • (ii)
    $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ for all $l\in\{1,\ldots,r\}$.
@@
As a consequence, there is a cycle of schools $(s_{1},\ldots,s_{r})$ in $G^{*}$ such that $(s_{1},s_{2})=(a,a^{\prime})$. Indeed, we can form this cycle departing from node $a$ in direction to node $a^{\prime}$ and moving around the graph, because after enter a node $s_{j}\neq s_{1}$ from an edge, we are always able to exit from another edge. Since $V^{*}$ is a finite set, we will eventually return to $s_{1}$. The construction of $E^{*}$ allow us to associated to $(s_{1},\ldots,s_{r})$ a group of students $i_{1},\ldots,i_{r}$ such that $i_{1}=i$, $\mu(i_{l})=s_{l}$ for all $l\in\{1,\ldots,r\}$, and $\mu^{\prime}(i_{l})=s_{l+1}=\mu(i_{l+1})$ for all $l\in\{1,\ldots,r\}$ [modulo $r$]. In particular, although $s_{1},\ldots,s_{r}$ are not necessarily different, we can assume that the students $i_{1},\ldots,i_{r}$ are.
Since $\mu^{\prime}(i_{l})=s_{l+1}$, when $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ the stability of $\mu$ ensures that $j\succ_{s_{l+1}}i_{l}$ for all $j\in\mu(s_{l+1})$. Hence, the stability of $\mu^{\prime}$ implies that any edge between $s_{l+1}$ and $s_{l+2}$ is associated to a student that improves we she moves from $\mu$ to $\mu^{\prime}$. In particular, $\mu^{\prime}(i_{l+1})P_{i_{l+1}}\mu(i_{l+1})$. Therefore, $\mu^{\prime}(i_{1})P_{i_{1}}\mu(i_{1})$ ensures that $\mu^{\prime}(i_{l})P_{i_{l}}\mu(i_{l})$ for all $l\in\{1,\ldots,r\}$.
We conclude that $(i_{1},\ldots,i_{r})$ satisfies properties (i) and (ii).∎